### v1

# Exact electromagnetic multipole expansion using elementary current multipoles

Abstract
Multipole expansion plays an important role in the description of electromagnetic scatterers, allowing them to be accurately characterized by a small set of expansion coefficients. However, to describe electromagnetic excitations inside a scatterer, the current density in it should be decomposed into current multipoles. Such current expansion includes nonradiating current configurations that are absent in the classical field-based expansion. Unfortunately, the use of current multipoles has so far been limited by the absence of an exact and general expression for the current multipole moments beyond their point-multipole approximation. Here, for the first time to our knowledge, we derive such an expression and present the exact mapping relations between the classical and current multipole moments. We use our theory to calculate the scattering and extinction cross sections for large, wavelength-scale, optical scatterers supporting multipole excitations up to the sixth order. Our results show perfect agreement with the Mie theory. The expressions are valid for electromagnetic scatterers of arbitrary sizes and shapes without restrictions on the multipole orders. They describe all possible current configurations, including those that do not produce any fields in the far-field zone. Current multipoles complement the existing theory for electromagnetic multipole expansion, while their minimalistic and universal character makes them a convenient tool for characterizing and designing diverse electromagnetic scattering systems of arbitrary complexity.
I Introduction
The radiation and scattering of electromagnetic fields are among the most fundamental phenomena in physics. Their rigorous theoretical treatment is essential in many areas of science and technology, from telecommunications [1] to atmospheric science [2, 3] and astrophysics [4]. Recently, scattering theory has become particularly relevant for the rapidly growing field of nanophotonics, in which the resonant optical response of nanostructured media has created new paradigms for optical field manipulation and enhanced light-matter interactions [5, 6]. For example, Mie multipole resonances [7, 8, 9, 10, 11] have been extensively explored for the realization of various optical phenomena, including Kerker effect [12, 13, 14], toroidal and anapole excitations [15, 16], and bound states in the continuum [17, 18, 19]. The engineering of multipole excitations has also been used in various functional structures, such as antireflection coatings [20], Huygens metasurfaces [21, 22], nanoscale lasers [23], metamaterials [24, 25], as well as systems exhibiting structural colors [26], collective resonances [27, 28, 29, 30], and enhanced nonlinear optical effects [31, 32].
Describing and designing resonant micro- and nanoscatterers can be done using electromagnetic multipole expansion. It allows one to characterize every finite-sized scattering system with a relatively small set of expansion coefficients corresponding to mutually orthogonal electromagnetic fields of different-order multipoles [33]. Since the original Mie solutions [34], multipole expansion has been approached in many ways [35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. However, regardless of the approach, the classical field-based expansion does not fully capture the actual current oscillations within the system. For example, electromagnetic anapoles are absent in this expansion, as they do not radiate into the far field [45, 46, 47, 48], so that their theoretical description requires the artificial separation of the toroidal parts of the multipole moments [49, 43, 11]. Moreover, if the scatterer is a part of a larger scattering system, its own scattered field is difficult to isolate. Therefore, the current oscillations within that scatterer cannot be easily determined through the classical field-based expansion [37, 50]. Finally, the multipole moments in the Cartesian field-based expansion [40] lack a generalized expression for an arbitrary order $l$, which limits their use to structures supporting only low-order multipole modes.
To overcome these limitations, a complementary multipole expansion has been proposed [37]. It expands the scattering current density
$\textbf{J}(\textbf{r})=-i\omega\epsilon_{0}[\epsilon_{\text{r}}(\mathbf{r})-\epsilon_{\text{s}}]\textbf{E}(\textbf{r})$ (1)
in terms of the current multipoles. In the point-multipole approximation, the moments of these multipoles are introduced by the following multipole tensor:
$\bm{M}^{(l)}_{\text{point}}=\frac{i}{(l-1)!\,\omega}\int\limits^{\infty}_{-\infty}\textbf{J}(\textbf{r})\textbf{r}^{l-1}d^{3}\textbf{r},$ (2)
similarly to the moments of the primitive Cartesian multipoles [51]. In the above equations, $\epsilon_{\text{r}}(\mathbf{r})$ is the position-dependent relative electric permittivity and $\epsilon_{\text{s}}$ is its value in the surrounding medium, $l$ is the multipole order, and $\mathbf{J}(\mathbf{r})\textbf{r}^{l-1}$ is the outer product of vector $\mathbf{J}(\mathbf{r})$ and l1l-1 vectors $\mathbf{r}$. The current multipoles are not divided into electric and magnetic multipoles and represent very simple configurations of linear current elements (see Fig. 1), which facilitates their use in designing individual and arrayed scatterers with prescribed scattering characteristics [52, 53, 48, 50, 54, 55].
Figure 1: Examples of elementary current configurations in the excitations of a current dipole ($l$ = 1), quadrupole ($l$ = 2), octupole ($l$ = 3), hexadecapole ($l$ = 4), and triacontadipole ($l$ = 5), with moments $p_{x}$, $Q_{yx}$, $O_{zxy}$, $H_{xxyz}$, and $T_{xxxyz}$, respectively. Current multipoles of high orders, such as 4 and 5, are relatively common, since often, they are constituents of classical multipoles of lower orders. In particular, current hexadecapoles are constituents of classical electric quadrupoles and magnetic octupoles, whereas current triacontadipoles appear in classical electric octupoles and magnetic hexadecapoles; see Eqs. (24)-(42). In numerical calculations presented in this work, we consider classical multipoles up to the electric and magnetic hexacontatetrapoles ($l$ = 6) [56] which include current multipoles of orders $l$ up to 8.
Unfortunately, Eq. (2) is not valid beyond the long-wavelength (point-multipole) approximation and cannot be used to evaluate the multipole moments and their contributions to the scattering and extinction cross sections for wavelength-scale and larger scatterers. So far, the only method for obtaining the exact current multipole moments for such scatterers was to calculate them from the exact moments of the classical electric and magnetic multipoles, i.e., from the multipole expansion coefficients $a_{\text{E}}(l,m)$ and $a_{\text{M}}(l,m)$, where $l$ and $m$ are the expansion orders [37, 50, 54, 55]. However, because current multipoles include nonradiating current configurations that are absent in the classical multipole expansion, the mapping from the classical to current multipole moments is underdetermined, so that some of the current multipole moments cannot be unambiguously evaluated. The lack of a complete theoretical framework for the current multipole expansion has so far prevented researchers from using the approach to its full potential, despite its clear advantages over the classical multipole expansion.
Here, we derive the long-sought general expression for the current multipole moments of any order $l$. The expression is valid for scatterers of arbitrary sizes and shapes. The obtained general current-multipole tensor reads:
$\bm{M}^{(l)}_{\text{exact}}=\frac{i}{\omega}\frac{(2l-1)!!}{(l-1)!}\int\limits_{-\infty}^{\infty}\textbf{J}(\textbf{r})\textbf{r}^{l-1}\frac{j_{l-1}(kr)}{(kr)^{l-1}}d^{3}\textbf{r},$ (3)
where $(2l-1)!!=(2l-1)\cdot(2l-3)\cdot...\cdot 5\cdot 3\cdot 1$ is the double factorial, $j_{l-1}(kr)$ are the spherical Bessel functions of the first kind, and $k$ is the wavenumber in the surrounding medium. The derivation of this expression is presented in Appendix A. Equation (3) allows the moments, i.e., the elements of the tensor $\bm{M}^{(l)}_{\text{exact}}$, to be calculated directly and unambiguously from the scattering current density distribution $\mathbf{J}(\mathbf{r})$ defined in Eq. (1). Several examples of these multipole moments are presented in Table 1, and the corresponding configurations of linear current elements in the point-multipole approximation are shown schematically in Fig. 1. Expressions in Table 1 illustrate the fact that every current multipole moment corresponds to a unique monomial containing a Cartesian vector component of the scattering current density ($J_{x}$, $J_{y}$, $J_{z}$) and, for $l\geq 2$, a product of normalized Cartesian coordinates ($x/r$, $y/r$, $z/r$). These moments can be called elementary as they constitute the simplest Cartesian configurations of oscillating currents. At the same time, they are exact and applicable to scatterers of arbitrary sizes and shapes, which we demonstrate in the next section.
Table 1: Exact expressions for the current multipole moments illustrated schematically in Fig. 1. The names of the multipoles follow the same convention as in Ref. [57].
$l$ Name Symbol Expression
1 Dipole $p_{x}$ $\frac{i}{\omega}\int\limits_{-\infty}^{\infty}j_{0}\left(kr\right)J_{x}(\textbf{r})\,d^{3}\textbf{r}$
2 Quadrupole $Q_{yx}$ $\frac{3i}{\omega k}\int\limits_{-\infty}^{\infty}j_{1}\left(kr\right)J_{y}(\textbf{r})\frac{x}{r}d^{3}\textbf{r}$
3 Octupole $O_{zxy}$ $\frac{15i}{2\omega k^{2}}\int\limits_{-\infty}^{\infty}j_{2}\left(kr\right)J_{z}(\textbf{r})\frac{xy}{r^{2}}d^{3}\textbf{r}$
4 Hexadecapole $H_{xxyz}$ $\frac{35i}{2\omega k^{3}}\int\limits_{-\infty}^{\infty}j_{3}\left(kr\right)J_{x}(\textbf{r})\frac{xyz}{r^{3}}d^{3}\textbf{r}$
5 Triacontadipole $T_{xxxyz}$ $\frac{315i}{8\omega k^{4}}\int\limits_{-\infty}^{\infty}j_{4}\left(kr\right)J_{x}(\textbf{r})\frac{x^{2}yz}{r^{4}}d^{3}\textbf{r}$
II Numerical verification
To verify the correctness of Eq. (3), we used it to numerically calculate the multipole contributions to the scattering and absorption cross sections for wavelength-scale scatterers across the visible and near-infrared spectral range. We did this by calculating the multipole expansion coefficients $a_{\text{E}}(l,m)$ and $a_{\text{M}}(l,m)$ in terms of the elements of tensor $\bm{M}^{(l)}_{\text{exact}}$ and then evaluated their contributions to the cross sections in question.
The exact expressions for coefficients $a_{\text{E}}(l,m)$ and $a_{\text{M}}(l,m)$ in terms of the generalized current multipole moments are derived in Ref. [58]. The moments of the two expansions are just simple linear combinations of each other. For example, the classical electric and magnetic dipole moments are written as
$a_{\text{E}}(1,0)=$ $\,\sqrt{2}C_{1}p_{z}+\frac{7\sqrt{2}}{3}C_{3}(3O_{xxz}+3O_{yyz}$ (4)
$-O_{zxx}-O_{zyy}+2O_{zzz}),$
$a_{\text{E}}(1,\pm 1)=$ $\,C_{1}(\mp p_{x}+ip_{y})\mp\frac{7}{3}C_{3}[2O_{xxx}-O_{xyy}$ (5)
$-O_{xzz}+3O_{yxy}+3O_{zxz}\mp i(3O_{xxy}$
$-O_{yxx}+2O_{yyy}-O_{yzz}+3O_{zyz})],$
$a_{\text{M}}(1,0)=$  $\,5\sqrt{2}iC_{2}(-Q_{xy}+Q_{yx}),$ (6)
$a_{\text{M}}(1,\pm 1)\!=$  $\,5C_{2}[-Q_{xz}+Q_{zx}\pm i(Q_{yz}-Q_{zy})],$ (7)
where $C_{1}=-ik^{3}/(6\pi\epsilon E_{0})$, $C_{2}=-k^{4}/(60\pi\epsilon E_{0})$, and $C_{3}=-ik^{5}/(210\pi\epsilon E_{0})$, with $E_{0}$ being the incident field amplitude and $\epsilon=\epsilon_{0}\epsilon_{\text{s}}$. The relations for a set of higher-order multipoles are presented in Appendix B. Additionally, in Ref. [58] we provide a Python code to automatically derive expressions for an arbitrary order $l$.
After obtaining $a_{\text{E}}(l,m)$ and $a_{\text{M}}(l,m)$, the multipole contributions to the normalized scattering and extinction cross sections at each order $l$ are given by
$Q_{\text{scat,E/M}}(l)=\frac{\pi}{k^{2}S}\sum_{m=-l}^{l}(2l+1)|a_{\text{E/M}}(l,m)|^{2},$ (8)
$Q_{\text{ext,E}}(l)=-\frac{\pi}{k^{2}S}\sum_{m=-1,+1}(2l+1)\operatorname{Re}[ma_{\text{E}}(l,m)],$ (9)
$Q_{\text{ext,M}}(l)=-\frac{\pi}{k^{2}S}\sum_{m=-1,+1}(2l+1)\operatorname{Re}[a_{\text{M}}(l,m)],$ (10)
where $S$ is the geometric cross section of the scatterer ($S=\pi R^{2}$ for a spherical nanoparticle of radius $R$). The corresponding multipole contributions to the normalized absorption cross sections are then calculated as
$Q_{\text{abs,E/M}}(l)=Q_{\text{ext,E/M}}(l)-Q_{\text{scat,E/M}}(l).$ (11)
In the numerical calculations, $\mathbf{J}(\mathbf{r})$ was obtained using Eq. (1) from the electric field distribution calculated in Comsol Multiphysics for a silicon sphere of diameter 600 nm (Fig. 2) and a silver sphere of diameter 400 nm (Fig. 3). In both cases, we assumed the scatterers to be embedded in PMMA, which is a dielectric material commonly used in nanofabrication. The optical constants of Si, Ag, and PMMA were taken from Refs. [59, 60, 61, 62]. The scatterers were illuminated by a plane wave linearly polarized along the $x$ axis and propagating along the $z$ axis, which is appropriate for the extinction cross sections defined by Eqs. (9) and (10). Figures 2 and 3 show the calculated cross sections. The points correspond to the approach based on the numerical evaluation of the current multipole moments using Eq. (3), while the lines show the direct predictions of the Mie theory. The approaches are seen to be in excellent agreement.
Note that Eqs. (4) and (5) for the classical electric dipole moments contain the current dipole and octupole moments, while Eqs. (6) and (7) for the classical magnetic dipole moments contain the current quadrupole moments. This shows that the magnetic dipoles in the field-based expansion and the electric quadrupoles in the current expansion are in fact the excitations of the same order. In general, every classical electric multipole moment of order $l$ can be expressed as a superposition of current multipole moments of orders $l$ and $l+2$, while every classical magnetic multipole moment can be written as a superposition of current multipole moments of order $l+1$.
Figure 2: Numerical verification of the exact expression for the current multipole moments [Eq. (3)] using the optical response of a silicon sphere of diameter 600 nm embedded in PMMA. The verification is done by evaluating the multipole contributions to the normalized scattering cross sections $Q_{\text{scat}}$ (a, b) and absorption cross sections $Q_{\text{abs}}=Q_{\text{ext}}-Q_{\text{scat}}$ (c, d) defined in Eqs. (8)-(11). The data points correspond to the contributions of moments $a_{\text{E}}(l,m)$ and $a_{\text{M}}(l,m)$ that are obtained by numerically calculating the current multipole moments $M^{(l)}_{\text{exact}}$ in Comsol Multiphysics with the help of Eqs. (1) and (3) and converting them to the classical multipole moments using Ref. [58] and Eqs. (4)-(7) and (24)-(42). The solid lines are the corresponding cross sections obtained from the Mie theory. The insets show the magnifications of the peaks corresponding to the electric and magnetic triacontadipoles ($l=5$) and hexacontatetrapoles ($l=6$). In the main plots, the curves for the subsequent multipoles are vertically shifted for clarity, and some of them are multiplied by a constant factor, as indicated. The results of numerical multipole expansion and Mie theory overlap perfectly (up to the numerical precision of Comsol), which proves the correctness of the expressions presented in this work.
Figure 3: Numerical verification of the current multipole expansion using the optical response of a silver sphere of diameter 400 nm embedded in PMMA. Similarly to Fig. 2, the solid lines represent the cross sections calculated using the Mie theory, while the points correspond to the cross sections obtained numerically in Comsol Multiphysics using the current multipole expansion (see Eqs. (1), (3), (4)-(7), (24)-(42), and Ref. [58]). In the studied spectral range, the scattering cross sections ($Q_{\text{scat}}$) have significant contributions of (a) the electric multipoles of order $l\leq 6$ (up to hexacontatetrapoles) and (b) the magnetic multipoles of order $l\leq 3$ (up to octupoles), while the main contributions to the absorption cross sections ($Q_{\text{abs}}$) come from the electric multipoles of $l\geq$4, as shown in (c). Similarly to Fig. 2, there is an excellent agreement between the numerical multipole expansion and Mie theory, except for the small deviations due to the finite numerical precision of the Comsol simulations.
Conclusions
In summary, we have introduced an expression for the general and exact current multipole tensor [see Eq. (3)], which can be used to completely characterize any localized current configurations, including those that do not radiate into the far field. The expressions for current multipole moments remain very simple also for large multipole orders (see Table 1). Furthermore, unlike the classical field-based multipoles, current multipoles are not divided into electric and magnetic types, which reflects their general, fundamental nature. Despite the simple and intuitive form of the current multipole moments, they can be used to quantitatively determine the multipole contributions to the scattering and extinction cross sections for scatterers of arbitrary sizes and shapes. We have verified our expression numerically for wavelength-scale spherical scatterers supporting higher-order multipole excitations and found an excellent agreement of our generalized expansion with the Mie theory. Our expansion simplifies the analysis of multipole excitations and the design of arbitrarily sized individual and arrayed electromagnetic scatterers, which makes it useful in the field of optical metamaterials and metasurfaces, as well as other electromagnetic scattering systems, from atoms and molecules to radio-frequency antennas.
The data that support the findings of this article are openly available [58].
The authors acknowledge the support of the Research Council of Finland (Grants No. 347449, 353758 and 368485). R. K. has received funding from the European Union’s Horizon Europe programme for research and innovation under the Marie Skłodowska-Curie Grant Agreement No. 101060306 (NExIA). For computational resources, the authors acknowledge the Aalto University School of Science “Science-IT” project and CSC – IT Center for Science, Finland.
Appendix A Derivation of Equation (3).
To expand the scattering current density J(r), one can use the identity
$\textbf{J}(\textbf{r})=\int\textbf{J}(\textbf{r}^{\prime})\delta(\textbf{r}^{\prime}-\textbf{r})d^{3}\textbf{r}^{\prime},$ (12)
where $\delta(\textbf{r}^{\prime}-\textbf{r})$ is the three-dimensional Dirac delta function, and $\textbf{r}^{\prime}$ stands for a coordinate inside the scatterer. Expanding the delta function into a Taylor series about $\textbf{r}^{\prime}=0$, we obtain
$\textbf{J}(\textbf{r})=\,i\omega\sum_{l=1}^{\infty}\sum_{\hat{\textbf{v}}=\hat{\textbf{x}},\hat{\textbf{y}},\hat{\textbf{z}}}\sum_{a=0}^{l-1}\sum_{b=0}^{l-a-1}M^{(l)}(\hat{\textbf{v}},a,b)\hat{\textbf{v}}\qquad\qquad$
$\times\frac{(-1)^{l}(l-1)!}{a!b![l-(a+b+1)]!}\frac{\mathrm{d}^{a}}{\mathrm{d}x^{a}}\frac{\mathrm{d}^{b}}{\mathrm{d}y^{b}}\frac{\mathrm{d}^{l-a-b-1}}{\mathrm{d}z^{l-a-b-1}}\delta(\textbf{r}),\qquad$ (13)
where $\hat{\textbf{x}}$, $\hat{\textbf{y}}$, and $\hat{\textbf{z}}$ are the unit vectors along the corresponding Cartesian coordinate axes and $M^{(l)}(\hat{\textbf{v}},a,b)$ are the elements of the multipole tensor $\bm{M}^{(l)}$ given in Eq. (2). These elements can be written as
$M^{(l)}(\hat{\textbf{v}},a,b)=\frac{i}{(l-1)!\,\omega}\int J_{v}(\textbf{r})x^{a}y^{b}z^{l-a-b-1}d^{3}\textbf{r},$ (14)
where $J_{v}(\textbf{r})$ is the v-component of vector $\textbf{J}(\textbf{r})$. Equation (14) coincides with Eq. (2) and is valid only if the size of the scatterer, $L$, satisfies the condition $kL\ll 1$.
A similar multipole expansion can be written for the vector potential A(r). Note that, in contrast to the electric and magnetic fields, the field of the vector potential is produced by electromagnetic anapoles as well. In the Lorentz gauge, the vector potential generated by a localized scattering current density obeys the Helmholtz equation:
$(\nabla^{2}+k^{2})\textbf{A}(\textbf{r})=-\mu\textbf{J}(\textbf{r}),$ (15)
where $k$ and $\mu$ are the wavenumber and magnetic permeability in the surrounding medium. The solution of this equation can be written as the following integral over the volume of the scatterer:
$\textbf{A}(\textbf{r})=\mu\int\textbf{J}(\textbf{r}^{\prime})G(\textbf{r}^{\prime}-\textbf{r})d^{3}\textbf{r}^{\prime},$ (16)
where $G(\textbf{r}^{\prime}-\textbf{r})=e^{ik|\textbf{r}^{\prime}-\textbf{r}|}/4\pi|\textbf{r}^{\prime}-\textbf{r}|$ is the Green’s function for a point source at a coordinate $\textbf{r}^{\prime}$ that is proportional to the spherical Hankel function of the first kind and zeroth order $h_{0}^{(1)}(k|\textbf{r}^{\prime}-\textbf{r}|)$. Using the identity $G(\textbf{r}^{\prime}-\textbf{r})=(ik/4\pi)h_{0}^{(1)}(k|\textbf{r}^{\prime}-\textbf{r}|)$ and expanding the Green’s function in Eq. (16) into a Taylor series about $\textbf{r}^{\prime}=0$, we obtain
$\textbf{A}(\textbf{r})=\,\frac{k^{3}}{4\pi\epsilon\omega}\sum_{l=1}^{\infty}\sum_{\hat{\textbf{v}}=\hat{\textbf{x}},\hat{\textbf{y}},\hat{\textbf{z}}}\sum_{a=0}^{l-1}\sum_{b=0}^{l-a-1}M^{(l)}(\hat{\textbf{v}},a,b)\hat{\textbf{v}}\qquad$
$\times\frac{(-1)^{l-1}(l-1)}{a!b![l-(a+b+1)]}\frac{\mathrm{d}^{a}}{\mathrm{d}x^{a}}\frac{\mathrm{d}^{b}}{\mathrm{d}y^{b}}\frac{\mathrm{d}^{l-a-b-1}}{\mathrm{d}z^{l-a-b-1}}h_{0}^{(1)}(kr),\qquad$ (17)
which is valid for $kL\ll 1$ and is the same as Eq. (32) in Ref. [37].
Equation (16) is general and can be used to find the exact expressions for the current multipole moments by expanding $G(\textbf{r}^{\prime}-\textbf{r})$ in terms of spherical harmonics $Y_{i,j}(\theta,\phi)$. This leads to the following well-known result [33]:
$\textbf{A}(\textbf{r})=$ $\,ik\mu\,\sum_{l=1}^{\infty}\,\sum_{\hat{\textbf{v}}=\hat{\textbf{x}}\,\hat{\textbf{y}},\hat{\textbf{z}}}\,\sum_{m=-l+1}^{l-1}h_{l-1}^{(1)}(kr)$ (18)
$\times Y_{l-1,m}(\theta,\phi)\tilde{M}^{(l)}(\hat{\textbf{v}},m)\hat{\textbf{v}}$,
where $r$, $\theta$, and $\phi$ are the spherical coordinates and $\tilde{M}^{(l)}(\hat{\textbf{v}},m)$ is defined as
$\tilde{M}^{(l)}(\hat{\textbf{v}},m)=\int J_{v}(\textbf{r})j_{l-1}(kr)Y_{l-1,m}^{*}(\theta,\phi)d^{3}\textbf{r}.$ (19)
Let us find several lowest-order coefficients $\tilde{M}^{(l)}(\hat{\textbf{v}},m)$ with the spherical harmonics written in Cartesian coordinates:
$\tilde{M}^{(1)}(\hat{\textbf{v}},0)=\frac{1}{2}\sqrt{\frac{1}{\pi}}\int J_{v}(\textbf{r})j_{0}(kr)d^{3}\textbf{r},$ (20)
$\tilde{M}^{(2)}(\hat{\textbf{v}},0)=\frac{1}{2}\sqrt{\frac{3}{\pi}}\int J_{v}(\textbf{r})z\frac{j_{1}(kr)}{r}d^{3}\textbf{r},$
$\tilde{M}^{(2)}(\hat{\textbf{v}},\pm 1)\!=\!\mp\frac{1}{2}\sqrt{\frac{3}{2\pi}}\int J_{v}(\textbf{r})(x\pm iy)\frac{j_{1}(kr)}{r}d^{3}\textbf{r},$
$\tilde{M}^{(3)}(\hat{\textbf{v}},0)=\frac{1}{4}\sqrt{\frac{5}{\pi}}\int J_{v}(\textbf{r})(3z^{2}-r^{2})\frac{j_{2}(kr)}{r^{2}}d^{3}\textbf{r},$
$\tilde{M}^{(3)}(\hat{\textbf{v}},\pm 1)=\mp\frac{1}{2}\sqrt{\frac{15}{2\pi}}\int J_{v}(\textbf{r})(xz\pm iyz)\frac{j_{2}(kr)}{r^{2}}d^{3}\textbf{r}.$
Each $\tilde{M}^{(l)}(\hat{\textbf{v}},m)$ can be written as a linear superposition of the following normalized multipole moments:
$M^{(l)}(\hat{\textbf{v}},a,b)=$ $\,\frac{i}{\omega}\frac{(2l-1)!!}{(l-1)!}\int J_{v}(\textbf{r})$ (21)
$\times x^{a}y^{b}z^{l-a-b-1}\frac{j_{l-1}(kr)}{(kr)^{l-1}}d^{3}\textbf{r}.\qquad$
These moments form the general current multipole tensor defined in Eq. (3). In the limit of $kr\ll 1$, $j_{l-1}(kr)/(kr)^{l-1}$ approaches 1/(2l1)!!1/(2l-1)!! [33], and Eq. (21) converges to Eq. (14). Correspondingly, Eq. (3) converges to Eq. (2).
The expression for the scattering current density in terms of the exact current multipole moments can be obtained from Eq. (12). Expanding the Dirac delta function in this equation in terms of spherical harmonics, we find:
$\textbf{J}(\textbf{r})=$ $\,\frac{k}{\pi}\sum_{l=1}^{\infty}\,\sum_{\hat{\textbf{v}}=\hat{\textbf{x}},\hat{\textbf{y}},\hat{\textbf{z}}}\,\sum_{m=-l+1}^{l-1}\frac{2l-1}{r^{2}}j_{l-1}(kr)$ (22)
$\times Y_{l-1,m}(\theta,\phi)\tilde{M}^{(l)}(\hat{\textbf{v}},m)\hat{\textbf{v}}$,
where each coefficient $\tilde{M}^{(l)}(\hat{\textbf{v}},m)$ is written in terms of multipole moments $M^{(l)}(\hat{\textbf{v}},a,b)$. The contribution of each current multipole to the near field of a scatterer can be assessed either from Eq. (18) or from Eq. (22), since the electric field is proportional to both A(r) and J(r). For example, if the characteristic size of the scatterer $L$ satisfies the condition $kL\ll 1$, the contributions of the current dipole, quadrupole, and octupole scale proportionally to $\textbf{{D}}/L$, $\textbf{{Q}}/L^{2}$, and $\textbf{{O}}/L^{3}$, respectively, since
$h_{0}^{(1)}(kL)=-ie^{ikL}\frac{1}{kL},$
$h_{1}^{(1)}(kL)=-e^{ikL}\frac{kL+i}{(kL)^{2}},$
$h_{2}^{(1)}(kL)=ie^{ikL}\frac{(kL)^{2}+3ikL-3}{(kL)^{3}}.$ (23)
Functions $j_{l-1}(kL)$ scale similarly. For larger scatterers, all the terms in the expressions for $h_{l-1}^{(1)}(kL)$ or $j_{l-1}(kL)$ must be included.
Appendix B Classical multipole moments as linear superpositions of current multipole moments.
Exact expressions for $a_{\text{E}}(l,m)$ and $a_{\text{M}}(l,m)$ with $l\geq 2$ in terms of current multipole moments up to triacontadipoles are given in Eqs. (24)-(42), supplementing the expressions for the electric and magnetic dipoles given in Eqs. (4)-(7). The factors $C_{1}$-$C_{3}$ in the expressions are defined below Eq. (7) and the two new factors are $C_{4}=-k^{6}/(350\pi\epsilon E_{0})$ and $C_{5}=-ik^{7}/(4410\pi\epsilon E_{0})$. The expressions for any order $l$ can be derived automatically, following instructions provided in Ref. [58].
$a_{\text{E}}(2,0)=$ $\,-\sqrt{6}\,C_{2}\,\left(\,Q_{xx}+Q_{yy}-2Q_{zz}\,\right)$ (24)
$-\sqrt{6}\,C_{4}\,(\,H_{xxxx}+H_{xxyy}-4H_{xxzz}+H_{yxxy}+H_{yyyy}$
$\qquad\qquad-4H_{yyzz}+3H_{zxxz}+3H_{zyyz}-2H_{zzzz}\,)$
$a_{\text{E}}(2,\pm 1)=$  $\,\mp\,3C_{2}\left(\,Q_{xz}+Q_{zx}\right)+\,3iC_{2}\left(\,Q_{yz}+Q_{zy}\,\right)$ (25)
 $\mp\,2C_{4}\,\left(\,4H_{xxxz}-H_{xyyz}-H_{xzzz}+5H_{yxyz}-H_{zxxx}-H_{zxyy}+4H_{zxzz}\,\right)$
 $-\,2iC_{4}\left(\,4H_{yyyz}-H_{yxxz}-H_{yzzz}+5H_{xxyz}-H_{zxxy}-H_{zyyy}+4H_{zyzz}\,\right)$
$a_{\text{E}}(2,\pm 2)=$  $\,3C_{2}\left[\,Q_{xx}-Q_{yy}\mp i\left(Q_{xy}+Q_{yx}\,\right)\,\right]$ (26)
$+\,C_{4}\,\left(\,3H_{xxxx}-7H_{xxyy}-2H_{xxzz}+7H_{yxxy}-3H_{yyyy}+2H_{yyzz}+5H_{zxxz}-5H_{zyyz}\,\right)$
 $\mp\,2iC_{4}\left(\,4H_{xxxy}-H_{xyyy}-H_{xyzz}-H_{yxxx}+4H_{yxyy}-H_{yxzz}+5H_{zxyz}\,\right)\,$
$a_{\text{E}}(3,0)=$  $\,2\sqrt{3}\,C_{3}\,\left(\,2O_{xxz}+2O_{yyz}+O_{zxx}+O_{zyy}-2O_{zzz}\,\right)$ (27)
+ $+\,2\sqrt{3}\,C_{5}\,(\,15T_{xxxxz}+15T_{xxyyz}-20T_{xxzzz}+15T_{yxxyz}+15T_{yyyyz}-20T_{yyzzz}$
 $\qquad\qquad-\,3T_{zxxxx}-6T_{zxxyy}+24T_{zxxzz}-3T_{zyyyy}+24T_{zyyzz}-8T_{zzzzz}\,)$
$a_{\text{E}}(3,\pm 1)=$ $\,\mp C_{3}\left(\,3O_{xxx}+O_{xyy}-4O_{xzz}+2O_{yxy}-8O_{zxz}\,\right)$ (28)
$+iC_{3}\left(\,3O_{yyy}+O_{yxx}-4O_{yzz}+2O_{xxy}-8O_{zyz}\,\right)$
 $\mp\,3C_{5}\,(\,4T_{xxxxx}+3T_{xxxyy}+5T_{yxyyy}-T_{xyyyy}+4T_{xzzzz}+3T_{xyyzz}+5T_{yxxxy}$
$\qquad\quad-27T_{xxxzz}-30T_{yxyzz}+15T_{zxxxz}+15T_{zxyyz}-20T_{zxzzz}\,)$
$+3iC_{5}\,(\,4T_{yyyyy}+3T_{yxxyy}+5T_{xxyyy}-T_{yxxxx}+4T_{yzzzz}+3T_{yxxzz}+5T_{xxxxy}$
$\qquad\quad-27T_{yyyzz}-30T_{xxyzz}+15T_{zxxyz}+15T_{zyyyz}-20T_{zyzzz}\,)$
$a_{\text{E}}(3,\pm 2)=$ $\,\sqrt{10}\,C_{3}\,[\,-2O_{xxz}+2O_{yyz}-O_{zxx}+O_{zyy}\pm 2i\left(\,O_{xyz}+O_{yxz}+O_{zxy}\right)\,]$ (29)
$-3\sqrt{10}\,C_{5}\,(\,5T_{xxxxz}-9T_{xxyyz}-2T_{xxzzz}+9T_{yxxyz}-5T_{yyyyz}+2T_{yyzzz}$
$\qquad\qquad\quad-T_{zxxxx}+6T_{zxxzz}+T_{zyyyy}-6T_{zyyzz}\,)$
[…]
$\qquad\qquad\quad-T_{zxxxy}-T_{zxyyy}+6T_{zxyzz}\,)$
$a_{\text{E}}(3,\pm 3)=$ $\,\sqrt{15}\,C_{3}\left[\,\pm\left(\,O_{xxx}-O_{xyy}-2O_{yxy}\,\right)-i\left(\,2O_{xxy}+O_{yxx}-O_{yyy}\,\right)\,\right]$ (30)
$\pm\sqrt{15}\,C_{5}\,(\,4T_{xxxxx}-21T_{xxxyy}-15T_{yxyyy}+13T_{yxxxy}+3T_{xyyyy}+3T_{xyyzz}$
$\qquad\qquad\quad-3T_{xxxzz}+6T_{yxyzz}+7T_{zxxxz}-21T_{zxyyz}\,)$
$+\sqrt{15}\,i\,C_{5}\,(\,4T_{yyyyy}-21T_{yxxyy}-15T_{xxxxy}+13T_{xxyyy}+3T_{yxxxx}+3T_{yxxzz}$
$\qquad\qquad\quad-3T_{yyyzz}+6T_{xxyzz}+7T_{zyyyz}-21T_{zxxyz}\,)\,]$
$a_{\text{M}}(2,0)=7\sqrt{6}\,i\,C_{3}\left(\,O_{xyz}-O_{yxz}\,\right)$ (31)
$a_{\text{M}}(2,\pm 1)=-7C_{3}\,[\,O_{xyy}-O_{xzz}-O_{yxy}+O_{zxz}\pm i\left(\,O_{xxy}-O_{yxx}+O_{yzz}-O_{zyz}\,\right)\,]$ (32)
$a_{\text{M}}(2,\pm 2)=7C_{3}\left[\,\mp\left(\,O_{xxz}-O_{yyz}-O_{zxx}+O_{zyy}\,\right)+i\left(\,O_{xyz}+O_{yxz}-2O_{zxy}\,\right)\,\right]$ (33)
$a_{\text{M}}(3,0)=\frac{5}{2}\sqrt{3}\,i\,C_{4}\,\left(\,-H_{xxxy}-H_{xyyy}+4H_{xyzz}+H_{yxxx}+H_{yxyy}-4H_{yxzz}\,\right)$ (34)
$a_{\text{M}}(3,\pm 1)=$ $\,-\frac{5}{4}C_{4}[\,H_{xxxz}+11H_{xyyz}-4H_{xzzz}-10H_{yxyz}-H_{zxxx}-H_{zxyy}+4H_{zxzz}$ (35)
$\qquad\quad\mp\,i\left(\,H_{yyyz}+11H_{yxxz}-4H_{yzzz}-10H_{xxyz}-H_{zyyy}-H_{zxxy}+4H_{zyzz}\,\right)\,]$
$a_{\text{M}}(3,\pm 2)=$ $\,\frac{5}{4}\sqrt{10}\,C_{4}\,[\,\pm\left(\,2H_{xxyy}-2H_{xxzz}-2H_{yxxy}+2H_{yyzz}+2H_{zxxz}-2H_{zyyz}\,\right)$ (36)
$\qquad\qquad\quad+i\left(H_{xxxy}-H_{xyyy}+2H_{xyzz}-H_{yxxx}+H_{yxyy}+2H_{yxzz}-4H_{zxyz}\,\right)\,]$
$a_{\text{M}}(3,\pm 3)=$ $\,\frac{5}{4}\sqrt{15}\,C_{4}\,[\,H_{xxxz}-H_{xyyz}-2H_{yxyz}-H_{zxxx}+3H_{zxyy}$ (37)
$\qquad\quad\quad\pm\,i\left(\,H_{yyyz}-H_{yxxz}-2H_{xxyz}-H_{zyyy}+3H_{zxxy}\,\right)\,]$
$a_{\text{M}}(4,0)=14\sqrt{5}\,i\,C_{5}\,\left(\,3T_{xxxyz}+3T_{xyyyz}-4T_{xyzzz}-3T_{yxxxz}-3T_{yxyyz}+4T_{yxzzz}\,\right)$ (38)
$a_{\text{M}}(4,\pm 1)=$ $\,-7C_{5}\,[\,3T_{xxxyy}-3T_{xxxzz}+3T_{xyyyy}-21T_{xyyzz}+4T_{xzzzz}-3T_{yxxxy}$ (39)
$\qquad\quad-3T_{yxyyy}+18T_{yxyzz}+3T_{zxxxz}+3T_{zxyyz}-4T_{zxzzz}$
$\qquad\quad\mp\,i\,(\,3T_{yxxyy}-3T_{xxxxy}+3T_{yxxxx}-21T_{yxxzz}+4T_{yzzzz}-3T_{xxyyy}$
$\qquad\qquad-3T_{yyyzz}+18T_{xxyzz}+3T_{zxxyz}+3T_{zyyyz}-4T_{zyzzz}\,)\,]$
$a_{\text{M}}(4,\pm 2)=$ $\,-7\sqrt{2}C_{5}\,[\,\pm\,(\,T_{xxxxz}+15T_{xxyyz}-6T_{xxzzz}-15T_{yxxyz}-T_{yyyyz}+6T_{yyzzz}$ (40)
$\qquad\qquad\quad-T_{zxxxx}+6T_{zxxzz}+T_{zyyyy}-6T_{zyyzz}\,)$
$\qquad\qquad+2i\,(\,3T_{xxxyz}-4T_{xyyyz}+3T_{xyzzz}-4T_{yxxxz}+3T_{yxyyz}$
$\qquad\qquad\quad+3T_{yxzzz}+T_{zxxxy}+T_{zxyyy}-6T_{zxyzz}\,)\,]$
$a_{\text{M}}(4,\pm 3)=$  $\,7\sqrt{7}C_{5}\,[\,3T_{xxxyy}-3T_{yxxxy}-T_{xyyyy}+T_{yxyyy}+3T_{xyyzz}$ (41)
$\qquad\qquad-3T_{xxxzz}+6T_{yxyzz}+3T_{zxxxz}-9T_{zxyyz}$
$\qquad\quad\pm\,i\,(\,3T_{yxxyy}-3T_{xxyyy}-T_{yxxxx}+T_{xxxxy}+3T_{yxxzz}$
$\qquad\qquad-3T_{yyyzz}+6T_{xxyzz}+3T_{zyyyz}-9T_{zxxyz}\,)\,]$
$a_{\text{M}}(4,\pm 4)=$ $\,-7\sqrt{14}C_{5}\,[\,\mp\left(\,T_{xxxxz}-3T_{xxyyz}-3T_{yxxyz}+T_{yyyyz}-T_{zxxxx}+6T_{zxxyy}-T_{zyyyy}\,\right)$ (42)
$\qquad\qquad\quad+i\,(\,3T_{xxxyz}-T_{xyyyz}+T_{yxxxz}-3T_{yxyyz}-4T_{zxxxy}+4T_{zxyyy}\,)\,]$